find the equation f(x) of degree three that has zeroes at 1 and 1+i such that f(1)=-8?
Answer by Edwin McCravy(20065)

This is a trick question because if f(x) has a zero at 1,
that means f(1) = 0.  So f(1) cannot equal -8.  There can
therefore be no such polynomial function.

I will therefore pretend there was a typo and that the f(1) 
should have been f(2) and that your question was this
legitimate question instead:

find the equation f(x) of degree three that has zeroes
at 1 and 1+i such that f(2)=-8?
The polynomial function that has zeros at r1, r2, r3, ..., rn 
is of the form

f(x) = k(x-r1)(x-r2)···(x-rn)

You are given zeros 1 and 1+i 

Therefore we have r1 = 1, r2 = 1+i, and r3 = 1-i 

[The reason for r3 = 1-i is the fact that if a polynomial 
with real coefficients has imaginary zero a+bi, it also has a zero
which is its conjugate a-bi.

So we have:

f(x) = k(x-1)[x-(1+i)]{x-(1-i)]

Remove the parentheses inside the brackets:

f(x) = k(x-1)[x-1-i]{x-1+i]

Put parentheses around the (x-1)'s:

f(x) = k(x-1)[(x-1)-i][(x-1)+i]

Multiply the bracketed expressions using FOIL:

f(x) = k(x-1)[(x-1)²+i(x-1)-i(x-1)-i²]

The two middle terms cancel:

f(x) = k(x-1)[(x-1)²-i²]

Replace i² by (-1) 

f(x) = k(x-1)[(x-1)²-(-1)]

f(x) = k(x-1)[(x-1)²+1]

f(x) = k(x-1)[x²-2x+1+1]

f(x) = k(x-1)(x²-2x+2)

f(x) = k(x³-2x²+2x-x²+2x-2)

f(x) = k(x³-2x²+2x-x²+2x-2)

f(x) = k(x³-3x²+4x-2)

Now since we are given that f(2) = -8,

we substitute x=2 

f(2) = k(2³-3·2²+4·2-2)

We substitute -8 for f(2)

-8 = k(8-3·4+8-2)

-8 = k(8-12+6)

-8 = k(2)

-8 = 2k

-4 = k

Substitute -4 for k in

f(x) = k(x³-3x²+4x-2)

f(x) = -4(x³-3x²+4x-2)

f(x) = -4x³+12x²-16x+8

Edwin